OpenStudy (anonymous):

Solve for the variable: -2v – 3 = 2

OpenStudy (jackellyn):

The purpose is to get the variable alone. So add 3 to both sides -2v -3+3 = 2+3 -2v = 5 Divide by -2 on both sides -2v/-2 =5/-2 v = -5/2

OpenStudy (anonymous):

-2v - 3 = 2 +3 -2v = 5/ -2 v = -2 1/2 or -5/2